OpenStudy (richyw):

Set up the double integral giving the volume of the solid bounded by \(y=x^2,\quad z=3y,\quad z=2+y\)

OpenStudy (richyw):

this is messing with my head. I am taking R to be \[\left[3y,\,2+y\right]\times\left[ 0, 1\right]\]

OpenStudy (richyw):

sorry that is z=3y to z=2+y and y=0 to y=1

OpenStudy (richyw):

It's been awhile since I have done this haha

OpenStudy (anonymous):